You are in charge of securing a rectangular work area by fencing it in. Two sides of the area measure 20 feet and two sides measure 10 feet. How many feet of fencing are needed?
Answer by stanbon(75887)

You are in charge of securing a rectangular work area by fencing it in. Two sides of the area measure 20 feet and two sides measure 10 feet. How many feet of fencing are needed?
-------
Ans: 2(20+10) = 2*30 = 60 ft
=================================
Cheers,
Stan H.
================